OpenStudy (anonymous):

Solve equation: log18(4v+9)=log15(5v-10). how would u solve this. please help

OpenStudy (anonymous):

are 18 and 15 the base or is it just the natural log?

OpenStudy (whpalmer4):

\[\log_{18}(4v+9) = \log_{15}(5v-10)\]Correct?

OpenStudy (anonymous):

yes whpalmer4

OpenStudy (anonymous):

and dreamer02

OpenStudy (whpalmer4):

Well, I've done it graphically, but haven't figured out how to get an exact solution (or if it can be done)...

OpenStudy (anonymous):

okay

OpenStudy (whpalmer4):

I wonder if perhaps noting that 18 and 15 have a GCF of 3 and using the change of base formula might get you anywhere...

OpenStudy (anonymous):

i wondering the same thing but im not sure

OpenStudy (whpalmer4):

Nah, it doesn't seem to do the job...and even Mathematica doesn't seem to be able to find an exact solution. I guess you have to do it numerically or graphically

OpenStudy (anonymous):

i would go with numerically

OpenStudy (anonymous):

because it want's it to be solved not graphed

OpenStudy (whpalmer4):

Then I guess you're trying to find the value of \(v\) such that \[\frac{\ln (9+4v)}{\ln 18} - \frac{\ln(5v-10)}{\ln 15} = 0\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

Look here http://www.wolframalpha.com/input/?i=%28log_%2818%29+%284v%2B9%29%29%3D%28log_%2815%29+%285v-10%29%29 I don't think there is an analytic way of solving it sorry...

OpenStudy (anonymous):

wait, i got it.

OpenStudy (anonymous):

okay i kinda get it

OpenStudy (anonymous):

what

OpenStudy (anonymous):

so, to solve the problem, multiply @whpalmer4 's equation by log(18)*log(15)

OpenStudy (anonymous):

the top part or everything

OpenStudy (anonymous):

no, even that way won't work...

OpenStudy (whpalmer4):

Not really necessary to do so, those are just constants, and they aren't going to bother your root finder...

OpenStudy (anonymous):

i was thinking about using ln(a)-ln(b)=ln(a/b)=0 and from there a/b=1, but you still have the constants...

OpenStudy (anonymous):

which then become powers... there just isn't a way of solving this analytically.

OpenStudy (anonymous):

yea it looks like there is no way hah